OpenStudy (anonymous):

perform the indicated operation and, if possible, simplify. Show all work. Write final answer exactly as a fraction in lowest terms. 3/14+2/5

OpenStudy (anonymous):

3/14 = 15/70 (multiply top and bottom by 5) 2/5 = 28/70 (multiply top and bottom by 14) now that the denominators are the same just add the tops so it equals 43/70